I tried a manual test of the new user space model. Since this is a user space process running it should have no affect on the rest of the front end system (which it didn't):
$ (cd /opt/rtcds/caltech/c1/target/c1dnn/c1dnnepics && ./startupC1)
$ taskset -c 6 /opt/rtcds/caltech/c1/target/c1dnn/bin/c1dnn -m c1dnn
Unfortunately, the process died with an "ADC TIMEOUT" error. I'm investigating why.
Once we confirm the model runs, we'll add the appropriate SHMEM IPC connections to connect it to the c1lsc model.
I tried moving the model to c1ioo, where there are plenty of free cores sitting idle, and the model seems runs fine. I think the problem was just CPU contention on the c1lsc machine, where there were only two free cores and the kernel was using both for all the rest of the normal user space processes.
So there are two options:
I'm going to start testing cpuset offline to figure out exactly what would need to be done.
Telescope front lens to wall distance 25 cm, GigE camera lenght 6 cm and cat6 cable 2cm
Atm3, Existing short camera can has 16cm lenght to lexan guard on viewport. Available 2" od periscope tube lenght is 8cm. The one in use 16 cm long.
Note: we can fabricate a lite cover with tube that would accomodate longer telescope.
Can we calibrate the AR coated M5018-SW and compare it's performance agains the 2" periscope
Look at the Edmond Optics 3" od camera lens with AR
Atm1, Now I can see dust. This is much better. The focus is not right yet.
Atm2, Chamber viewport wiped and image refocused. Actually I was focusing on the dust.
We don't really have to calibrate the lens, just the CCD, which we've done. It's more about knowing the true aperture size to know how much solid angle you're capturing to infer the total amount of scatter. For our custom lens tubes this is the ID of the retaining ring.
The Edmund Optics lens tube looks tempting, but itcomes at a price. Thorlabs sells lens tubes that offer a more flexibility than what we have right now, so I bought a few different ones, and also more 150mm 2" lenses. This will allow for more compact solutions and offer some in-situ focusing ability that doesn't require detaching the lens tube like now. Should be here in a couple of days, then we'll be able to enclose the GigE camera in the viewport can with a similar field of view we have now.
I also bought a collimation package for the AS port fiber stuff so we can move ahead with the ringdown measurements and also mode spectroscopy.
I was looking at the ASDC channel on dataviewer, and toggling various settings like whitening gain. At some point, the signal just froze. So I quit dataviewer and tried restarting it, at which point it complained about not being able to connect to FB. This is when I brought up the CDS_OVERVIEW medm screen, and noticed the frozen 1pps indicator lights. There was certainly something going on with the end FEs, because I was able to ping the machine, but not ssh into it. Once the 1pps lights came back, I was able to ssh into c1iscex and c1iscey, no problems.
Could it be that some of the mx processes stalled, but the systemctl routine automatically restarted them after some time?
An mx_stream glitch would have interrupted data flowing from the front end to the DAQ, but it wouldn't have affected the heartbeat. The heartbeat stop could mean either that the front end process froze, or the EPICS communication stopped. The fact that everything came back fine after a couple of minutes indicates to me that the front end processes all kept running fine. If they hadn't I'm sure the machines would have locked up. The fact that you couldn't connect to the FE machine is also suspicious.
My best guess is that there was a network glitch on the martian network. I don't know how to account for the fact that pings still worked, though.
Could it be that some of the mx processes stalled, but the systemctl routine automatically restarted them after some time?
So this wasn't just an EPICS freeze? I don't see how this had anything to do with any of the work I did earlier today. I didn't modify any of the running front ends, didn't touch either of the end station machines or the DAQ, and didn't modify the network in any way. I didn't leave anything running.
If you couldn't access test points then it sounds like it was more than just EPICS. It sounds like maybe the end machines somehow fell of the network momentarily. Was there anything else going on at the time?
While working on the IFO tonight, I noticed that the blinky status lights on c1iscex and c1iscey were frozen (but those on the other 3 FEs seemed fine). But all other lights on the CDS overview screen were green I couldn't access testpoints from these machines, and the EPICS readbacks for models on these FEs (e.g. Oplev servo inputs outputs etc) were frozen at some fixed value. This lasted for a good 5 minutes at least. But the blinky lights started blinking again without me doing anything. Not sure what to make of this. I am also not sure how to diagnose this problem, as trending the slow EPICS records of the CPU execution cycle time (for example) doesn't show any irregularity.
We took a closer look at the AS55 demod board today. The procedure was to just be as thorough as possible, and check the behaviour of the circuit (both Transfer Function and Noise) stage by stage. Checking the transfer function was the key.
During this process, we found that the reason why the Q channels had lower noise than the I channels was because of the gain of the AD829 stage of the circuit was 0dB rather than 4dB (which is what it should be according to the component values used). Specifically, resistor R12, which is supposed to be 1.30kohm, was measured to be 1.03kohm. Replacing this resistor, the transfer functions (see Attachment #1) and noise levels (see Attachment #2) match the expectations from LISO. Some notes:
Assming the whitened ADC noise level is much below this (should only be ~10nV/rtHz), and given the measured sensing element of 6.2e8 V/m, this means that the dark noise sets a maximum achievable sensitivity of 2e-16m/rtHz.
To figure out what (if anything) is to be done next, we need to first figure out what is the goal. In the end, we care about DARM and not MICH. The optical gain for the former is ~300x the latter, so the dark noise contribution gets scaled by this factor (giving us a number of 7e-19 m/rtHz). There are certainly many noises above that level which have to be handled first. Indeed, looking at the DARM spectrum from DRFPMI lock back in March 2016, it looks like the current 1f DRMI (with coils de-whitened) Michelson sensitivity is within a factor of 2 of DARM in the full lock (albeit with vertex DoFs on 3f signals, and no coil de-whitening). Koji pointed out that we need to consider the photodiode resonant circuit itself too.
TODO: Upload all this onto the DCC
I've been preparing for testing Gabriele's deep neural network MICH/PRCL reconstruction. No changes to the front end have been made yet, this is all just prep/testing work.
We have been unable to get Gabriele's nn.c code running in kernel space for reasons unknown (see tests described in previous post). However, Rolf recently added functionality to the RCG that allows front end models to be run in user space, without needing to be loaded into the kernel. Surprisingly, this seems to work very well, and is much more stable for the overall system (starting/stopping the user space models will not ever crash the front end machine). The nn.c code has been running fine on a test machine in this configuration. The RCG version that supports user space models is not that much newer than what the 40m is running now, so we should be able to run user space models on the existing system without upgrading anything at the 40m. Again, I've tested this on a test machine and it seems to work fine.
The new RCG with user space support compiles and installs both kernel and user-space versions of the model.
Koji suggested looking at the output of the AS55 demod board on a fast oscilloscope to look for differences in the two channel outputs (if there is some high-frequency oscillations, for example, we could miss this information in the SR785 spectra). Besides, I was only looking at spectra out to a few kHz on the SR785. I grabbed this data with a 300MHz BW Tektronix oscilloscope (battery mode) today. Input impedance of both channels were set to 1Mohm, and the measurement was made with the RFPD input terminated, output of the daughter board is what is measured. The vertical scaling of the channels was set to the minimum allowed, 1mV/div.
Attachment #1 shows that there is indeed a visible difference between the two channels - the (noisier) I channel has a much larger DC offset of ~5mV compared to the Q channel (I tried switching channels on the O'scope and the larger DC offset remained on the I channel, so seems real). There is also some kind of oscillation going on in the I channel, although the frequency is pretty low, with the peaks spaced ~50us apart. Indeed, in the ASD of the acquired data, the excess power in the I channel at 20kHz and higher harmonics are evident (see Attachment #2). Anyway all of this points to something being anomalous on the daughter board I channel signal path - I will pull it out and monitor the outputs at various points along the signal path with the fast scope to see if I can narrow down what's going on where.
Both channels should be identical - I don't understand why the I channels are noisier than their Q counterparts. This is almost certainly a problem on the daughter board, as the orange traces are pretty much identical for both channels.
Wall StripTool traces showed that IMC has not been locked for at least 8 hours when I came in this morning. Going to the IMC autolocker log, it looks like the last timestamp was at ~6pm yesterday. Megatron was responding to ping, but I couldn't ssh into it. So I went over to the machine and did a hard-reboot via front panel power switch. The computer took ~10mins to come back online and respond to ping. Once it did, I was able to ssh into it. However, trying the usual commands to restart the IMC autolocker and FSS Slow loops didn't work. Specifically, monitoring the logfile with tail -f Autolocker.log, I would see that the autolocker seemed to get stuck after starting the "blinky" script. Trying to restart the process using sudo initctl restart MCautolocker, init would print to shell that the restart had worked, and reported the PID, but the logfile wouldn't update "live" as it should when tail is used with the -f option. All very strange .
Anyways, as a last resort, I kill -9'ed the PID for the init instance, and init automatically restarted the Autolocker - this did the trick, IMC is locked now and logfile seems to be getting updated normally.
I also cleared a bunch of matlab crash dump files in the home directory.
Attachment #1 - Measured / modelled noises for AS55 demod board. I've plotted quadrature sum of the LISO trace with the SR785 noise floor with input terminated to ground via 50ohm. Note that these measurements were made after all the changes in the marked up schematic in the previous elog were implemented.
The dark red curves were measured by shorting the inputs to D040179 to ground via 50ohms using some Pomona minigrabbers - I wanted to avoid ripping the daughter board out, but this probably explains the excess noise compared to the green trace at low frequencies. All other measurements were made with the board installed in the LSC rack eurocrate, with the LO input driven at the nominal level (I didn't measure this yesterday but a measurement from ~6months ago says that this level is 1.5dBm).
Steve reported problems getting the X arm locked. Alignment sliders were inaccessible. Eurocrate key turning reboots for c1susaux, c1auxex,c1auxey, c1iscaux and c1aux. Usual precautions were taken for ITMX.
This is becoming a once-a-week thing .
Here is the marked up schematic with the board as it is stuffed. Annoyingly, there is a capacitor (C1) which according to the schematic is supposed to be open, but is stuffed in our board. I can't find any elog about this, and its a pain to measure the value of this capacitance. I will upload all of this + LISO + noise model/measurements to a 40m AS55 daughter board DCC page.
This lower priced 1" apeture Navitar lens can be an option too.
Atm1, Now I can see dust. This is much better. The focus is not right yet.
I calculated a better lens solution for the ETMX side view with the simple python script that's attached. The camera is still not as close to the viewport as we would like, and now the front lens is almost all the up to the end of the tube. With a little more playing around there maybe a better way, especially if we expand the repertoire of focal lengths. Using Steve's wonderful camera fixture I put the beam spot in focus. I turned the camera sideways for better use of the field of view, and now the beam spot actually fills the center area of the beam, to the point where we probably don't want more magnification or else we start losing the tails of the Gaussian.
We'll take a serious of images tomorrow, and will have an estimate of the scatter loss by the end of tomorrow.
I worked on the daughter board a little more in the evening. I have somehow managed to make the dark noise ~25% worse [Attachment #1].
After making all these changes, I re-installed the card in the eurocrate and repeated the measurement. The Q channel noise was close to the expected value (~50nV/rtHz), but the I channel is twice as noisy. I will continue this investigation tomorrow.
I tried replacing the AD797s on the daughter board with OP27s, and saw no significant improvement in the electronics noise of the demod board. Note that according to LISO, in this configuration, the voltage noise of the Op27 is expected to dominate the total noise of the daughter board. Measurement condition was that the RFPD input was terminated, but the LO input was still being driven (SR785 input range is -50dBVpk for all traces, and the input ranging was set to "UpOnly"). Need to do a more systematic investigation to figure out where this excess noise is coming from. I will upload photos of the board later.
This supports the hypothesis that something is wonky on the daughter board, because the purple trace should only be the quad sum of the orange and green traces. I will pull it out and have a look.
these are not the SR785 settings that you're looking for
Gautam and Steve,
All 3 show the same noise level ~80 nV / rt Hz at 1 kHz as shown. Batteries ordered to be replaced in the top 2
We'll do more measurement to see how can we get to 4 nV / rt Hz specification level.
To get low noise measurements on the SR785, you have to have the input range set to -50 dB, not +20 dB. Its not within the powers of commercial electronics ADCs to give you a 10 nV noise floor with +10 V input signals. The SR560 has an input referred noise of 5 nV/rHz, so the output noise should be 5e-9 x 500 = 2.5 uV/rHz. Your picture shows it giving 1 uV RMS, so you also need to use the PSD units.
I keep adding traces to this plot, here is the most complete one I have now. Looks like the input noise to the D040179 (measured at "Q out" SMA jack of D990511 with RFPD input terminated) is ~10nV/rtHz. This supports the hypothesis that something is wonky on the daughter board, because the purple trace should only be the quad sum of the orange and green traces. I will pull it out and have a look.
Some other follow-ups on the questions raised at the meeting:
I measured the output voltage noise of the Q output of the AS55 Demod Board with the PSL shutter closed, using the SR785 (see Attachment #1). The measured noise is consistent with the expected number of ~120nV/rtHz around 100Hz. I had measured the gain of this board from RFPD input to Q output to be ~5.1: so if the PD dark noise is 16nV/rtHz, this would be amplified to ~80nV/rtHz. Still a discrepancy of ~50%. I didn't measure the noise with the PD input terminated. Added the noise of the demod board output with the RFPD input terminated. The level of ~100nV/rtHz seems consistent with the actual PD dark noise being ~80nV/rtHz, as their quadrature sum is around 130nV/rtHz. Need to dig up the schematics for the demod board + daughter board, and check against LISO, to see if this is consistent with what is expected.
Also - I think I was using the wrong value of the DC power on the AS55 photodiode for shot noise calculations - 13mW was for REFL55, not AS55. I did a crude measurement of the power by sticking the Ophir power meter (filter removed) in front of the AS55 PD with the Michelson flashing around, and noticed the maximum value registered was ~1.2mW. So in the DRMI lock, there would be ~2.4mW, which is 10x lower than the value I was assuming. I've made the correction in the NB code, for the next time the plot is generated. A more rigorous measurement would involve sticking the Ophir in front of the AS110 PD during a DRMI lock. The light from the AS port is split by a 50-50 BS to the AS55 and AS110 PDs (so measuring at AS110 is a reasonable proxy for power at AS55), and the AS110 signals are not used for triggering in the DRMI lock, so this is feasible.
how about calibrate the DC channels so that you can just get the acutal power levels from the trend data?
My last characterization of the AS55 PD was on Feb 2013. ELOG 8100
There I said the dark noise at the PD output was 16nV/rtHz. I don't have the measurement of the Voltage noise at the output of the demod board.
Note that the PD can only be limited by shot noise when the DC current is larger then 4mA.
I spent this weekend doing a more careful investigation of the DRMI noise. I think I have some new information/insights. Attachment #1 is the noise budget (png attached because pdf takes forever to upload, probably some ImageMagick problem. The last attachment is a tarball of the PDF). Long elog, so here are the Highlights:
The DAC noise is not limiting us anywhere when the coil de-whitening is switched on.
So I don't understand the measured Dark Noise level, and it is limiting us at frequencies > 200Hz. Some busted electronics in the input signal chain? Or can the LSC demod daughter board gain of ~5 explain the observed noise?
Edit 1730 9 Oct: I had missed out the factor of 5 gain in the demod board in calculating the shot noise curve. Attachment #7 shows the corrected shot noise level. Explicitly:
, where is to convert shot noise in W to displacement units.
This seems to be limiting us from saturating the dark noise once the coil de-whitening is engaged. But lack of coherence means the mechanism is not re-injection of SRCL/PRCL sensing noise? Need to think about what this means / how we can mitigate it.
I've also made several changes to the NB code - will push to git once I finish cleaning stuff up, but it is now much faster to make these plots and see what's what.
There are no more double sided tape on this table.
Gabriele had prepared a C code implementation of his NN for MICH/PRCL state estimation. He had been trying to get it going on some of the machines in WB, but was unsuccessful. The version of RCG he was trying to compile and run the code on was rather dated so we decided to give it a whirl on our new RCG3.4 here at the 40m. Just noting down stuff we tried here:
All models have been reverted to their state prior to this test, and everything on the CDS_OVERVIEW MEDM screen is green now.
Looks to have worked this time around.
controls@fb1:~ 0$ sudo dd if=/dev/sda of=/dev/sdc bs=64K conv=noerror,sync
33554416+0 records in
33554416+0 records out
2199022206976 bytes (2.2 TB) copied, 55910.3 s, 39.3 MB/s
You have new mail in /var/mail/controls
I was able to mount all the partitions on the cloned disk. Will now try booting from this disk on the spare machine I am testing in the office area now. That'd be a "real" test of if this backup is useful in the event of a disk failure.
The 4TB HGST drives have arrived. I've started the FB1 dd backup process. Should take a day or so.
Steve, can you please connect this fan to the rack power and remove this extra power supply?
Re-arranged the DC bench supply on the shelf in the PSL enclosure, whose only purpose seems to be to supply 12V to a fan attached to the rear of the PSL NPRO controller. Seems to be a waste of space! The fan was momentarily disconnected but has since been reconnected and is spinning again.
I've located the Stanford Research FS725 Rb reference unit. The question is where to put it. This afternoon Steve and I put it inside the little electronics rack next to 1X3, but in hindsight, this probably isn't such a great place for a timing reference as there are a bunch of Sorensen power supplies in there (and presumably the accompanying harmonics from these switching supplies).
The unit itself was repaired in 2015, and powering it on, it locked to the internal reference within a few minutes as prescribed in the manual.
MC Autolocker was umnhappy because c1iool0 was unresponsive and hence it couldn't write to the "C1:IOO-MC_AUTOLOCK_BEAT" channel. I keyed the crate and IMC locked almost immediately. I'm moving this channel into the RTCDS model as we did for the IFO_STATE EPICS channel so that the autolocker isn't dependant on c1iool0 (which was the whole point of migrating the IFO-STATE variable anyways). I also commented out all of these channels in /cvs/cds/caltech/target/c1iool0/autolocker.db so that there aren't duplicate channels.
Eurocrate key turning reboots for c1susaux, c1auxex,c1auxey, c1iscaux, c1iscaux2 and c1aux. Usual precautions were taken for ITMX. Did burtrestore for c1iscaux andc1iscaux2 in order to restore the LSC PD whitening gains.
Un-related to this work: input pointing into PMC was tweaked as the PMC_REFL spot was pretty bright.
In the end I decided to access the available spare DAC channels via the C1ASS model - for this purpose, I added a namespace block "TEST" in the C1ASS simulink model, which is a SISO block. Inside is just a single CDS filter module. My idea is to use the EXC of this filter module to inject excitations for measuring various couplings. Rather than have a simple testpoint, we also have the option of adding in some filter shapes in the filter module which could possibly allow a more direct read-off of some coupling TF. Recompiling the model went smooth - there was a crash earlier in the day which required me to hard-reboot c1lsc (and also restart all models on c1sus and c1ioo but no reboots necessary for those machines).
Note that to get the newly added channels to show up in the channel lists in DTT/AWGGUI etc, you need to ssh into fb1 and restart the daqd processes via sudo systemctl restart daqd_*. If I remember right, it used to be enough to do telnet fb 8088 followed by shutdown. This is no longer sufficient.
It took me a while to get the DRMI locking going again. The model restarts earlier in the evening had changed a bunch of EPICS channel settings (and out config scripts don't catch all of these settings). In particular, I forgot to re-enable the x3 digital gain for the ITMs, BS and SRM (necessitated by removing an analog x3 gain on the de-whitening boards). I was hesitant to spend time re-adjusting all damping / oplev loop gains because if we change the series resistor on the coil driver board, we will have to do this again. I also didn't want this arbitrary FM to be enabled in the SDF safe.snap. But maybe it's worth doing it anyways - if nothing it'll be good practise.
Once I hunted down all the setting diffs and tweaked alignment, the DRMI locks were pretty robust.
I had hoped to make some of these TF measurements tonight. But I realized I needed to look up a bunch of stuff in manuals/datasheets, and think about these measurements a little. I wasn't sure if the DW/AI board could drive a signal over 40m of BNC cabling so I added an SR560 (DC coupled, gain=1, low noise mode, 50ohm output used) to buffer the output. The Marconi's external modulation input is high impedance (100k) but for the AOM driver we want 50ohm. For the Marconi, the external input accepts 1Vrms max, while for the AOM driver, we want to drive a signal between 0V and 1V at most.
The general measurement setup is schematically shown in Fig 1. Questions to address:
Am I missing something?
There are at least 5 free DAC channels (4 if you discount the one channel from these that I am hijacking) available in the 1Y2 electronics rack.
Jamie's nice wiring diagram shows the topology - the actual DAC card sits in 1Y3 inside the c1lsc expansion chassis (while the c1lsc frontend itself is in 1X4). The output of the DAC goes via SCSI to an interface box (D080303) and then to some dewhitening/AI boards (D000316). There are a total of 16 DAC channels available, out of which 8 are used for the TTs, 2 are used for the DAFI model, and one is labeleld "From c1ioo 1X2" (I don't know what this one is for). So I'm going to use some of these channels for measuring the coupling of oscillator noise and intensity noise to MICH in the DRMI lock.
The de-whitening/AI board seems to be old - it has 2x 800Hz Butterworth LPFs and no notch for the clock frequency, but maybe this doesn't matter for the tests I have in mind. The AI board available on 1X2 is more modern but routing the DAC channels from 1Y2 to it is going to be some work.
I'm going to add my testpoint to c1daf given that it seems to be the least critical model on c1lsc.
EDIT: testpoints added to c1daf don't show up in the list of available channels - there was some issue with this model while we were getting the new RTCDS going. So I'm moving my temporary testpoint to c1cal instead.
Disclaimer: Wrong calibration factors! See https://nodus.ligo.caltech.edu:8081/40m/13391
The factors were indeed enormously off. The correct table reads:
I did subtract a 'dark' frame from the images, though not in the sense of your point 1, just an exposure of identical duration with the laser turned off. This was mostly to reduce the effect of residual light, but given similar initial conditions would somewhat compensate for the offset that pre-existing charge and electronics noise put on the pixel values. The white field is of course a difference story.
I wonder how close we can get to a white field by putting a thin piece of paper in front of the camera without lenses and illuminate it from the other side. A problem is of course the coherence if we use a laser source... Or we scrap any sort of screen/paper and illuminate directly with a strongly divergent beam? Then there wouldn't be a specular pattern.
I'm not sure I understand your point about the 1.5V/A. Just to make sure we're talking about the same thing I made a crude drawing:
The PD sees plenty of light at all times, and the 1.5V/uW came from a comparative measurement PD<-->Ophir (which took the place of the CCD) while adjusting the power deflected with the AOM, so it doesn't have immediate connection to the conversion gain of silicon in this case. I can't remember the gain setting of the PD, but I believe it was 0dB, 20dB at most.
GV Oct 6: This coupling is probably not correct - Finesse outputs TF magnitude in units of W/W, and not W/RIN.
Since I was foiled (by lack of DAC) in my attempt to measure the coupling of laser intensity noise to MICH in the DRMI (no arms) configuration, I decided to try understanding the effect with a simulation.
For this purpose, I used my DRMI Finesse model - this had mirror positions tuned for locking and photodiode demod phases tuned to give a sensing matrix model that wasn't too far from an actual measurement (within factor of a few). So the model seems okay for a first pass at estimating this coupling.
Measuring transfer functions in Finesse is straightforward - use the fsig command to modulate some quantity (in this case the input beam intensity), and use the pd2 detector to demodulate the effect of this modulation at the port of interest (in this case AS55_Q).
**Note that to apply a modulation to an input beam (i.e. Laser) in Finesse, the keyword for the "type" argument given to fsig is "amp" and not "amplitude" as the manual would had me believe. In fact, there seem to be quite a few such caveats. The best way to figure this out is to go to the pykat installation directory, find the file components.py, and look for the fsig_name for the component of interest. It is also indicated in the same file, via the canFsig argument, if that property of the component can be modulated for transfer function measurements.
Attachment #1 shows the result of such a sweep.
To estimate what the actual contribution to the displacement noise is, I used the DQ-ed MC transmission (recorded at 1024Hz) from the DRMI lock, computed the ASD using scipy.signal.welch, divided by the nominal MC transmission of ~15,000 counts to convert to RIN/rtHz. The RIN was then multiplied by the above calculated coupling function, and divided by the sensing matrix element for AS55_Q (in units of W/m) to give the curve shown in Attachment #2. If we believe the simulation, then Laser Intensity Noise shouldn't be the limiting noise between 10Hz-1kHz.
I will of course measure the actual coupling and see how it lines up with Attachment #1 - would be a nice additional validation of the Finesse model. I will also try using the Finesse model to estimate some other coupling transfer functions (e.g. Laser Frequency Noise, Oscillator Noise).
The absence of evidence is not evidence of absence.
Going through some astronomy CCD calibration resources (-), I gather that there are in general 3 distinct types of correction that are applied:
The flat-field calibration seems to be the most complicated - the idea is to use a source of known radiance, and capture an image of this known radiance with the CCD. Then assuming we know the source radiance well enough, we can use some math to back out what the actual response function of individual pixels are. Then, for an actual image, we would divide by this response-map to get the actual image. There are a number of assumptions that go into this, such as:
I am not sure what error is incurred by ignoring 2 and 3 in the list at the beginning of this elog, perhaps this won't affect our ability to estimate the scattered power from the test-masses to within a factor of 2. But it may be worth it to do these additional calibration steps.
I also wonder what the uncertainty in the 1.5V/A number for the photodiode is (i.e. how much do we trust the Ophir power meter at low power levels?). The datasheet for the PDA100A says the transimpedance gain at 60dB gain is 1.5 MV/A (into high impedance load), and the Si responsivity at 1064nm is ~0.25A/W, so naively I would expect 0.375 V/uW which is ~factor of 4 lower. Is there a reason to trust one method over the other?
Also, are the calibration factor units correct? Jigyasa reported something like 0.5nW s / ct in her report.
The incident power can be calculated as Pin =CF*Total(Counts-DarkCounts)/ExposureTime.
This did the trick - I simply ran
sudo systemctl restart daqd_*
on FB1, and now all the CDS overview lights are green again.
I thought I had done this already, but I realize that I was supposed to restart the daqd processes on FB1 (which is where they are running) and not on c1ioo .
Thanks Jamie for the speedy resolution!
From page 21 of T1100625, DAQ status "0x2000" means that the channel list is out of sync between the front end and the daqd. This usually happens when you add channels to the model and don't restart the daqd processes, which sounds like it might be applicable here.
It looks like open-mx is loaded fine (via "rtcds lsmod"), even though the systemd unit is complaining. I think this is because the open-mx service is old style and is not intended for module loading/unloading with the new style systemd stuff.
I was trying to set up a DAC channel to interface with the AOM driver on the PSL table.
[4072817.132040] c1als: Failed to allocate DAC channel.
[4072817.132040] c1als: DAC local 0 global 16 channel 4 is already allocated.
[4072817.132040] c1als: Failed to allocate DAC channel.
[4072817.132040] c1als: DAC local 0 global 16 channel 5 is already allocated.
[4072817.132040] c1als: Failed to allocate DAC channel.
[4072817.132040] c1als: DAC local 0 global 16 channel 6 is already allocated.
[4072817.132040] c1als: Failed to allocate DAC channel.
[4072817.132040] c1als: DAC local 0 global 16 channel 7 is already allocated.
[4073325.317369] c1als: Setting stop_working_threads to 1
controls@c1ioo:~ 130$ sudo systemctl status mx
● open-mx.service - LSB: starts Open-MX driver
Loaded: loaded (/etc/init.d/open-mx)
Active: failed (Result: exit-code) since Tue 2017-10-03 00:27:32 UTC; 34min ago
Process: 29572 ExecStop=/etc/init.d/open-mx stop (code=exited, status=1/FAILURE)
Process: 32507 ExecStart=/etc/init.d/open-mx start (code=exited, status=1/FAILURE)
Oct 03 00:27:32 c1ioo systemd: Starting LSB: starts Open-MX driver...
Oct 03 00:27:32 c1ioo open-mx: Loading Open-MX driver (with ifnames=eth1 )
Oct 03 00:27:32 c1ioo open-mx: insmod: ERROR: could not insert module /opt/3.2.88-csp/open-mx-1.5.4/modules/3.2.88-csp/open-mx.ko: File exists
Oct 03 00:27:32 c1ioo systemd: open-mx.service: control process exited, code=exited status=1
Oct 03 00:27:32 c1ioo systemd: Failed to start LSB: starts Open-MX driver.
Oct 03 00:27:32 c1ioo systemd: Unit open-mx.service entered failed state.
The two acA640-120gm Basler GigE-Cams have been calibrated. I used the collimated output of a fiber that carried the auxiliary laser light from the PSL table. With a non-polarizing beam splitter some of the light was picked off onto a PD, and I modified the RF amplitude of the AOM drive signal to vary the power coming out of the fiber. The fiber output was directed at a white paper, which was placed 1.06m from the front of the lens tube assembly, which is where the focal plane is. Using the Pylon Viewer App I made sure that the entirety of the beam spot was imaged onto the CCD. Since the camera sensor is 1/4" across, I removed the camera from the lens tube and instead placed the Ophir power meter head at the position of the sensor and measured the power reported versus PD voltage, which turned out to be 1.5 V/uW.
The camera was put back in place and I used the Pypylon package Gautam had stumbled upon to sweep the exposure time from 100us to 10ms at different light power settings including no laser light at all for background subtraction, and rather than keeping the full bitmap data for O(100s) of images I recorded only the quantities
I performed this procedure for both the 152 and 153 cameras and plotted the pixel sum and the pixel max vs the exposure time. All the exposures were taken at a gain setting of 100, which is the smallest possible setting (out of 100-600). To obtain the calibration factor I use the input power Pin=75nW in the 'safe' region 1ms to 10ms where the pixel sum looks smooth and the CCD is reportedly not saturated.
BS connections and damping restored.
The first Faraday isolater rejected beam path from the NPRO is fixed.
The FB1 dd backup process seems to have finished too - but I got the following message:
dd: error writing ‘/dev/sdc’: No space left on device
30523666+0 records in
30523665+0 records out
2000398934016 bytes (2.0 TB) copied, 50865.1 s, 39.3 MB/s
Running lsblk shows the following:
controls@fb1:~ 32$ lsblk
NAME MAJ:MIN RM SIZE RO TYPE MOUNTPOINT
sdb 8:16 0 23.5T 0 disk
└─sdb1 8:17 0 23.5T 0 part /frames
sda 8:0 0 2T 0 disk
├─sda1 8:1 0 476M 0 part /boot
├─sda2 8:2 0 18.6G 0 part /var
├─sda3 8:3 0 8.4G 0 part [SWAP]
└─sda4 8:4 0 2T 0 part /
sdc 8:32 0 1.8T 0 disk
├─sdc1 8:33 0 476M 0 part
├─sdc2 8:34 0 18.6G 0 part
├─sdc3 8:35 0 8.4G 0 part
└─sdc4 8:36 0 1.8T 0 part
While I am able to mount /dev/sdc1, I can't mount /dev/sdc4, for which I get the error message
controls@fb1:~ 0$ sudo mount /dev/sdc4 /mnt/HGSTbackup/
mount: wrong fs type, bad option, bad superblock on /dev/sdc4,
missing codepage or helper program, or other error
In some cases useful info is found in syslog - try
dmesg | tail or so.
Looking at dmesg, it looks like this error is related to the fact that we are trying to clone a 2TB disk onto a 1.8TB disk - it complains about block size exceeding device size.
sudo dd if=/dev/sda of=/dev/sdb bs=64K conv=noerror,sync
in a tmux session on nodus (as I did for chiara and FB1, latter backup is still running).
I've modified the __init.py__ file located at /ligo/apps/linux-x86_64/cdsutils-480/lib/python2.7/site-packages/cdsutils/__init__.py so that you can now simply import pyawg from cdsutils. On the control room workstations, iPython is set up such that cdsutils is automatically imported as "cds". Now this import also includes the pyawg stuff. So to use some pyawg function, you would just do (for example):
One could also explicitly do the import if cdsutils isn't automatically imported:
from cdsutils import awg
Linking this useful instructional elog from Chris here: https://nodus.ligo.caltech.edu:8081/Cryo_Lab/1748
? Why aren't you able to just import 'awg' directly? You shouldn't have to import it through cdsutils. Something must be funny with the config.
As Rana pointed out to me, one important fact to keep in mind w.r.t. DAC noise is that it can be non-linear. So the RMS of the DAC noise in a higher frequency band (say 60-100Hz) can be affected by the RMS of the requested DAC signal in some lower frequency band (say 10-20Hz). One test to see if this hypothesis can explain the difference @100Hz between the ITMX channels and BS channels I observed a couple of days ago is to see if the noise around 100Hz becomes lower when I enable a 20-40Hz bandstop in the digital signal chain.
The nodus backup too is now complete - however, I am unable to mount the backup disk anywhere. I tried on a couple of different machines (optimus, chiara and pianosa), but always get the same error:
mount: unknown filesystem type 'LVM2_member'
The disk itself is being recognized, and I can see the partitions when I run lsblk, but I can't get the disk to actually mount.
Doing a web-search, I came across a few blog posts that look like the problem can be resolved using the vgchange utility - but I am not sure what exactly this does so I am holding off on trying.
To clarify, I performed the cloning by running
After consulting with Jamie, we reached the conclusion that the reason why the root of FB1 is so huge is because of the way the RAID for /frames is setup. Based on my googling, I couldn't find a way to exclude the nfs stuff while doing a backup using dd, which isn't all that surprising because dd is supposed to make an exact replica of the disk being cloned, including any empty space. So we don't have that flexibility with dd. The advantage of using dd is that if it works, we have a plug-and-play clone of the boot disk and root filesystem which we can use in the event of a hard-disk failure.
This is not quite right. First of all, /frames is not NFS. It's a mount of a local filesystem that happens to be on a RAID. Second, the frames RAID is mounted at /frames. If you do a dd of the underlying block device (in this case /dev/sda*, you're not going to copy anything that's mounted on top of it.
What i was saying about /frames is that I believe there is data in the underlying directory /frames that the frames RAID is mounted on top of. In order to not get that in the copy of /dev/sda4 you would need to unmount the frames RAID from /frames, and delete everything from the /frames directory. This would not harm the frames RAID at all.
But it doesn't really matter because the backup disk has space to cover the whole thing so just don't worry about it. Just dd /dev/sda to the backup disk and you'll just be copying the root filesystem, which is what we want.